### v2

# Spectral Analysis and Decay Mechanisms of $1^{-+}$ Hybrid States in Light Meson Sector

Abstract
The exploration of exotic mesons, which transcend the conventional quark-antiquark framework, is pivotal for advancing our understanding of QCD and the strong interaction. Among these, states possessing the quantum numbers $J^{PC}=1^{-+}$, such as $\pi_{1}(1600)$, $\pi_{1}(2015)$, and the recently discovered $\eta_{1}(1855)$, have attracted significant attention due to their potential hybrid nature, which involves gluonic excitations. However, the physical interpretation of these states remains contentious, primarily due to inconsistencies between theoretical predictions and experimental observations regarding their decay widths and mass spectra. To address these challenges, we employ a potential model inspired by SU$(3)$ lattice gauge theory to calculate the masses of light-flavor $1^{-+}$ hybrid states and utilize the constituent gluon model to analyze their strong decay properties. Our mass calculations suggest that while the $\pi_{1}(1600)$ and $\eta_{1}(1855)$ masses are in close agreement with the predicted hybrid states, the corresponding decay widths for $\pi_{1}(1600)$ and $\pi_{1}(2015)$ do not support their classification as hybrids, implying alternative interpretations, such as tetraquark or molecular configurations. Conversely, the $\eta_{1}(1855)$ is consistent with a mixed hybrid state composed of $(u\bar{u}+d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$ components, with mixing angles ranging from $21^{\circ}$ to $74^{\circ}$. Additionally, we predict the masses and decay widths of yet-unobserved isoscalar and excited hybrid states, providing valuable targets for future experimental searches. This study not only clarifies the hybrid nature of certain exotic mesons but also contributes to the systematic classification of hybrid states within the meson nonet, thereby enhancing our comprehension of exotic hadronic states.
pacs:
12.39.Jh, 14.40.Rt
I Introduction
Approximately five decades ago, physicists formulated Quantum Chromodynamics (QCD) to describe the strong interaction tHooft:1972tcz ; Fritzsch:1973pi ; Gross:1973id ; Politzer:1973fx ; Gross:2022hyw . This development has significantly advanced our understanding of most hadrons and established the framework for hadron spectroscopy. However, our understanding of the gluon field’s role in the low-energy regime of QCD remains limited. Consequently, the study of conventional hadrons, where mesons consist of a quark and an antiquark, and baryons are composed of three quarks, appears insufficient. In response, the investigation of hybrid states, where gluons are incorporated as dynamic degrees of freedom, is crucial. This research is guided by two fundamental principles of QCD namely, asymptotic freedom Gross:1973id ; Politzer:1973fx and local color conservation Wilson:1974sk and is essential for a deeper understanding of non-perturbative QCD phenomena Fritzsch:1973pi ; Freund:1975pn .
The exploration of exotic mesons, which go beyond the conventional quark-antiquark framework, is crucial for advancing our understanding of QCD and the strong interaction. These mesons, characterized by unconventional quantum numbers, challenge our current theories and provide a unique opportunity to study the role of gluonic excitations. Among them, states with $J^{PC}=1^{-+}$, such as $\pi_{1}(1600)$ E852:1998mbq , $\pi_{1}(2015)$ E852:2004gpn and the recently observed $\eta_{1}(1855)$ BESIII:2022riz ; BESIII:2022iwi , have drawn significant attention due to their potential hybrid nature. However, interpreting the properties of these states remains a contentious issue, largely due to discrepancies between theoretical predictions and experimental observations in terms of their decay widths and mass spectra.
Several collaborations IHEP-Brussels-LosAlamos-AnnecyLAPP:1988iqi ; Aoyagi:1993kn ; E852:1997gvf ; VES:2001rwn ; CrystalBarrel:1998cfz have reported the observation of $\pi_{1}(1400)$, but its existence as a resonance remains controversial, as discussed in Meyer:2010ku . The $\pi_{1}(1600)$ has also been observed in various decay channels, including $\eta^{\prime}\pi$ Baker:2003jh , $\rho\pi$ Khokhlov:2000tk , $f_{1}(1285)\pi$ COMPASS:2009xrl , and $b_{1}(1235)\pi$ CLEO:2011upl . Over the years, many theorists have interpreted $\pi_{1}(1600)$ as a hybrid Narison:1999hg ; Shastry:2022mhk ; Chen:2023ukh ; Iddir:1988jd ; Benhamida:2019nfx ; Eshraim:2020ucw , though this paper presents an alternative viewpoint. Additionally, $\pi_{1}(1600)$ has been considered a candidate for a tetraquark state Chen:2008qw , given that gluons in a hybrid state can easily split into a quark-antiquark pair ($q\bar{q}$). Furthermore, some theories suggest that $\pi_{1}(1600)$ may be a mixing state of a molecular state and a tetraquark or hybrid Narison:2009vj . The $\pi_{1}(2015)$, observed exclusively in $b_{1}(1235)\pi$ E852:2004rfa and $f_{1}(1285)\pi$ E852:2004gpn by the BNL E852 experiment, remains poorly understood due to a lack of comprehensive experimental data. Lattice QCD calculations of radial excited states identify $\pi_{1}(2015)$ as a promising candidate for the first excited state of a hybrid Dudek:2010wm , though this contradicts the conclusions of the present work.
The $\eta_{1}(1855)$, discovered through partial-wave analysis of $J/\psi\to\gamma\eta_{1}(1855)\to\gamma\eta\eta^{\prime}$ BESIII:2022riz ; BESIII:2022iwi , is the first isoscalar particle with quantum numbers $J^{PC}=1^{-+}$. This discovery offers a unique opportunity to study exotic states. Extensive theoretical investigations have been carried out on the $\eta_{1}(1855)$. Within the framework of QCD sum rules, it has been interpreted as a four-quark state Wan:2022xkx . Additionally, the $\eta_{1}(1855)$ has been studied as a molecular state Dong:2022cuw ; Yang:2022rck ; Yan:2023vbh and as a hybrid state candidate using flux tube models Qiu:2022ktc , QCD sum rules Chen:2022qpd , and effective Lagrangian techniques Shastry:2022mhk . Like $\pi_{1}(1600)$, the physical nature of $\eta_{1}(1855)$ remains unresolved.
The limited number of discovered exotic particles and the scarcity of available data present significant challenges in accurately classifying these states. To address these challenges, we adopt a potential model inspired by SU(3) lattice gauge theory, which has demonstrated success in capturing gluonic contributions to hadron interactions, to calculate the masses of light-flavor $1^{-+}$ hybrid mesons. Additionally, the constituent gluon model is employed to analyze their strong decay properties, offering a robust framework to investigate the interplay of gluonic and quark-antiquark dynamics in hybrid mesons. We also predict the masses and decay widths of yet-undiscovered isoscalar and excited hybrid states, which provide valuable targets for experimental exploration. This study not only clarifies the hybrid nature of specific exotic mesons but also enhances the systematic classification of hybrid states within the meson nonet. Furthermore, conventional mesons are known to follow Regge trajectories Anisovich:2000kxa . Investigating whether hybrid states also adhere to this pattern is crucial for establishing a complete SU(3) multiplet, as discussed by other authors Chen:2023ukh ; Yan:2023vbh ; Qiu:2022ktc . By addressing these unresolved issues, our work contributes to a deeper understanding of the internal structure of exotic hadronic states and the dynamics of the strong interaction.
This paper is organized as follows: In Section II, we introduce the model for calculating the masses of $1^{-+}$ hybrid states and present theoretical masses for various flavor compositions. Section III describes the constituent gluon model used to calculate the decay widths of hybrid states. In Section IV, we discuss the decay widths of both discovered and undiscovered $1^{-+}$ particles, along with their dominant decay channels. Finally, Section V concludes with a discussion and summary.
II Masses of $1^{-+}$ hybrid
The quark potential model has proven highly effective in accurately predicting the masses of conventional mesons and baryons, prompting its application to hybrid systems. Recent advancements in lattice QCD have enabled the modeling of effective potentials for excited gluon fields, which facilitate interactions between quarks UKQCD:1998zbe ; Karbstein:2018mzo ; Capitani:2018rox ; Schlosser:2021wnr . These developments provide a foundation for the computation of hybrid meson masses within the $1^{-+}$ sector Chen:2023ukh . The effective potential for the lowest hybrid state with $J^{PC}=1^{-+}$ in the light-flavor sector is modeled as follows:
$V_{q\bar{q}g}(r)=\frac{A_{1}}{r}+A_{2}r^{2}+V_{0}+\xi_{2}\sqrt{\frac{b}{\xi_{1}}},$ (1)
where $q$ and $\bar{q}$ denote the light ($u$, $d$) and strange ($s$) quarks. The parameters are set as $A_{1}=0.0958$ GeV, $A_{2}=0.01035$ GeV3, $b=0.165$ GeV2, $\xi_{1}=0.04749$, and $\xi_{2}=0.5385$. The constant $V_{0}$ varies with flavor: $V_{0}^{(q\bar{q})}=-0.48$ GeV, $V_{0}^{(q\bar{s})}=-0.40$ GeV, and $V_{0}^{(s\bar{s})}=-0.33$ GeV.
We solve the spinless Salpeter equation:
$H=\sqrt{p_{q}^{2}+m_{q}^{2}}+\sqrt{p_{\bar{q}}^{2}+m_{\bar{q}}^{2}}+V_{q\bar{q}g}(r),$ (2)
where $p_{q}$ and $p_{\bar{q}}$ are the momenta of the quark and antiquark, respectively, and $m_{q}$, $m_{\bar{q}}$ their masses.
The wave functions are approximated using the Simple Harmonic Oscillator (SHO) basis:
$\psi_{nl}(r,\beta)=\beta^{3/2}\sqrt{\frac{2(2n-1)!}{\Gamma(n+l+\frac{1}{2})}}(\beta r)^{l}e^{-\frac{\beta^{2}r^{2}}{2}}L^{l+1/2}_{n-1}(\beta^{2}r^{2}),$ (3)
where $L^{\alpha}_{n}$ are the associated Laguerre polynomials, and $\beta$ is the oscillator parameter specific to each flavor configuration: $\beta^{H}_{q\bar{q}}=0.264~{}\text{GeV},\beta^{H}_{q\bar{s}}=0.271~{}\text{GeV}$ and $\beta^{H}_{s\bar{s}}=0.277~{}\text{GeV}$.
By solving the expectation value of energy, the corresponding mass ${M}_{nl}$ is obtained as
$M_{nl}(\beta)\equiv E_{nl}(\beta)=\langle\psi_{nl}|H|\psi_{nl}\rangle.$ (4)
Table 1: The masses of hybrid states with the radial quantum number $n$. All masses are in units of GeV.
$n$ 1 2 3 4 5
$q\bar{q}g$ 1.668 2.193 2.648 3.061 3.444
$q\bar{s}g$ 1.851 2.355 2.799 3.204 3.582
$s\bar{s}g$ 2.022 2.507 2.94 3.337 3.709
In this study, we use a potential model derived from lattice QCD to calculate the masses of light-flavor $1^{-+}$ hybrid states. The effective potential includes contributions from Coulomb-like terms, confinement effects, and gluonic excitations. The spinless Salpeter equation is solved using a harmonic oscillator wavefunction basis, allowing us to obtain the energy eigenvalues. The computed masses are presented in Table 1 and show close agreement with experimental data for $\pi_{1}(1600)$ and $\eta_{1}(1855)$, which correspond to $q\bar{q}g$ and $q\bar{s}g$ configurations, respectively. However, deviations are observed for $\pi_{1}(2015)$, suggesting it may belong to a different excitation class.
Additionally, we analyze the mass spectrum using Regge trajectories (see Fig. 1), which describe the linear relationship between squared masses of mesons and their quantum numbers. The observed adherence of hybrid mesons to Regge trajectories, similar to conventional mesons, suggests that hybrid states may exhibit regularities comparable to traditional meson families. This finding suggests that hybrids may exhibit regularities akin to those of ordinary mesons, paving the way for the systematic classification of hybrid mesons as part of a broader meson family.
The fitting result equation corresponding to Fig. 1 :
$M_{s\bar{s}g}^{2}=$ $M_{s0}^{2}+(n-1)\mu_{1},$
$M_{q\bar{q}g}^{2}=$ $M_{q0}^{2}+(n-1)\mu_{2},$ (5)
where $M_{s0}^{2}$ and $M_{q0}^{2}$ represent the ground state masses of $s\bar{s}g$ and $q\bar{q}g$, respectively.
The Reggie trajectory exhibited by mesons in Ref. Anisovich:2000kxa is
$M_{q\bar{q}}^{2}=M_{0}^{2}+(n-1)\mu,$ (6)
where $\mu_{1}=2.41$, $\mu_{2}=2.27$ and $\mu=1.25$.
In isospin partner mesons, the $s\bar{s}g$ and $n\bar{n}g$ flavor configurations often mix, with $n\bar{n}g$ denoting $(u\bar{u}+d\bar{d})/\sqrt{2}g$. The calculated mass of $s\bar{s}g$ is slightly higher than that of $\eta_{1}(1855)$, while the mass of $n\bar{n}g$ is slightly lower. This suggests that the $\eta_{1}(1855)$ state could be a hybrid of these two configurations. In Section IV, we analyze the strong decay behavior of the $\eta_{1}(1855)$, treating it as a mixture of $s\bar{s}g$ and $n\bar{n}g$.
Figure 1: (color online). The analysis of the ($n,M^{2}$) trajectories for the $1^{-+}$ states. The trajectory slopes are 2.41 and 2.27 for the $s\bar{s}g$ and $n\bar{n}g$ states, respectively. Here,the solid circles with gray denote calculated values of phenomenological theory and circles denote the theoretical values of Reggie trajectory, while the error bar correspond to the experimental data listed in Particle Data Grou (PDG) ParticleDataGroup:2024cfk .
III Decay model of the hybrid state
The decay properties of hybrid states are essential for understanding their structure and classification. Several theoretical models have been developed to investigate the decay widths of hybrid states, including the flux tube model Page:1998gz ; Isgur:1985vy ; Close:1994hc ; Barnes:1995hc , the constituent gluon model LeYaouanc:1984gh ; Iddir:1988jd ; Ishida:1991mx ; Kalashnikova:1993xb ; Swanson:1997wy ; Iddir:2000yb ; Ding:2006ya ; Iddir:2007dq ; Benhamida:2019nfx ; Farina:2020slb ; Tanimoto:1982eh ; Tanimoto:1982wy , QCD sum rules DeViron:1984svx ; Zhu:1998sv ; Zhang:2002id ; Chen:2010ic ; Huang:2010dc ; Huang:2016upt , and lattice QCD McNeile:2002az ; McNeile:2006bz ; Woss:2020ayi . In this work, we adopt the constituent gluon model, where the hybrid meson is represented as a quark-antiquark pair coupled to a transverse electric (TE) gluon ($J_{g}^{PC}=1^{+-}$) Chen:2023ukh . This gluon acts as a dynamic degree of freedom and can annihilate into a new quark-antiquark pair, which then combines with the original pair to form two mesons. This mechanism describes how hybrid states decay into two mesons, as illustrated in Fig. 2.
Figure 2: The mechanism of a hybrid state A decaying into B and C mesons through the constituent gluon model.
For the representation of the hybrid states the following notations are used:
[…]
$L_{q\bar{q}}$  : the relative orbital momentum between q and $\bar{q}$;
$S_{q\bar{q}}$  : the total quarks spin;
$J_{g}$  : the total gluon angular momentum;
[…]
Considering the gluon moving in the framework of the $q\overline{q}$ pair, the Parity of the hybrid will be Iddir:2007dq :
$P=\left(-\right)^{L_{q\bar{q}}+L_{g}}.$ (8)
The charge conjugation of hybrid states is expressed as:
$C=\left(-\right)^{L_{q\bar{q}}+S_{q\bar{q}}+1}.$ (9)
$S_{q\bar{q}}$ can be either 0 or 1, Eq. 8 and Eq. 9 imply that both $L_{q\bar{q}}+L_{\text{g}}$ and $L_{q\bar{q}}+S_{q\bar{q}}$ must be odd for $J^{PC}=1^{-+}$. Under the constraints of the Clebsch-Gordan coefficients in Eq. III, the lightest states of $J^{PC}=1^{-+}$ hybrid is presented in Table 2, which is categorized into two general classes: $L_{q\bar{q}}=0$ and $L_{g}=1$ which is usually referred as the gluon-excited hybrid (GE hybrid), and $L_{q\bar{q}}=1$ and $L_{g}=0$ which is usually referred as the quark-excited hybrid (QE hybrid).
Within the degree of freedom of the transverse electric (TE) gluon, the $1^{-+}$ hybrid state corresponds to the gluon-excited state (GE hybrid) LeYaouanc:1984gh .
Table 2: The quantum numbers of $J^{PC}=1^{-+}$ $\,q\bar{q}g$ hybrid mesons
$P$ $C$ $L_{q\bar{q}}$ $L_{\text{g }}$ $J_{g}$ $S_{q\bar{q}}$ $L$ $J$
-- $+$ 0 1 0 1 0 1
-- $+$ 0 1 1 1 1 1
-- $+$ 0 1 2 1 2 1
-- $+$ 1 0 1 0 1 1
In QCD theory, the Hamiltonian describing the transition of a hybrid state into two conventional mesons is as follows:
${H}_{I}=g\int\textup{d}^{3}\vec{x}~{}\bar{\psi}(\vec{x})\gamma_{\mu}\frac{\lambda^{a}}{2}\psi(\vec{x})A_{a}^{\mu}(\vec{x}).$ (10)
The operator relevant to the decay can be expressed by operators that encompass the annihilation of gluons and the creation of quark-antiquark pairs:
$H_{I}=g\sum_{s,s^{\prime},\lambda,c,c^{\prime},c_{g}}\int\frac{d^{3}\mathbf{p}\;d^{3}\mathbf{p}^{\prime}d^{3}\mathbf{k}}{\sqrt{2\omega_{g}}(2\pi)^{9}}(2\pi)^{3}\delta^{3}(\mathbf{p}-\mathbf{p^{\prime}}-\mathbf{k})$
×$\times\bar{u}_{\mathbf{p}sc}\gamma_{\mu}\frac{\lambda^{c_{g}}_{c,c^{\prime}}}{2}v_{-\mathbf{p}^{\prime}s^{\prime}c^{\prime}}b^{\dagger}_{\mathbf{p}sc}d^{\dagger}_{-\mathbf{p}^{\prime}s^{\prime}c^{\prime}}a^{c_{g}}_{\mathbf{k}\lambda}\varepsilon^{\mu}(\mathbf{k},\lambda).$ (11)
The flavor index of the quark, $c_{g}=1,2,\cdot\cdot\cdot 8$, has been omitted and the creation and annihilation operators satisfy the relationship:
$\{b_{\mathbf{p}sc},b^{\dagger}_{\mathbf{p}^{\prime}s^{\prime}c^{\prime}}\}$ $=$ $\{d_{\mathbf{p}sc},d^{\dagger}_{\mathbf{p}^{\prime}s^{\prime}c^{\prime}}\}=(2\pi)^{3}\delta^{3}(\mathbf{p}-\mathbf{p}^{\prime})\delta_{ss^{\prime}}\delta_{cc^{\prime}}$
$[a^{c_{g}}_{\mathbf{k}\lambda},a^{{c^{\prime}_{g}}^{\dagger}}_{\mathbf{k}^{\prime}\lambda^{\prime}}]$ $=$ $(2\pi)^{3}\delta^{3}(\mathbf{k}-\mathbf{k}^{\prime})\delta^{c_{g},c^{\prime}_{g}}\delta_{\lambda\lambda^{\prime}}$ (12)
The non-relativistic approximation function for hybrid and mesons are described as follows :
$|A(L_{g},L_{q\bar{q}},S_{q\bar{q}},J_{A},M_{J_{A}})\rangle=\sum\int\frac{d^{3}\mathbf{p}_{1}d^{3}\mathbf{p}_{2}d^{3}\mathbf{k}}{(2\pi)^{9}}(2\pi)^{3}\chi^{\mu}_{ss^{\prime}}\frac{\lambda^{c_{g}}_{c_{q},c_{\bar{q}}}}{4}\psi_{L_{q\bar{q}}M_{L_{q\bar{q}}}}\left(\frac{m_{\bar{q}}\mathbf{p}_{1}-m_{q}\mathbf{p}_{2}}{m_{q}+m_{\bar{q}}}\right)\;\psi_{L_{g}M_{L_{g}}}\left(\frac{\mathbf{k}(m_{\bar{q}}+m_{\bar{q}})-(\mathbf{p}_{1}+\mathbf{p}_{2})\omega_{g}}{(m_{q}+m_{\bar{q}})\omega_{g}}\right)\;$
$\langle L_{g},M_{L_{g}};1,\lambda_{g}|J_{g},M_{J_{g}}\rangle\langle L_{q\bar{q}},M_{L_{q\bar{q}}};J_{g},M_{J_{g}}|L,m^{\prime}\rangle\langle L,m^{\prime};S_{q\bar{q}},M_{S_{q\bar{q}}}|J_{A},M_{J_{A}}\rangle\delta^{3}(\mathbf{p}_{1}+\mathbf{p}_{2}+\mathbf{k}-\mathbf{P}_{A})b^{\dagger}_{\mathbf{p}_{1}s_{1}c_{1}}d^{\dagger}_{\mathbf{p}_{2}s_{2}c_{2}}a^{{c_{g}}^{\dagger}}_{\mathbf{k}\lambda_{g}}|0\rangle.$ (13)
The final $B$ meson’s state is given by:
$|B$ $(L_{B},S_{B},J_{B},M_{J_{B}})\rangle=\sum_{M_{L_{B}},M_{S_{B}}}\int\frac{d^{3}\mathbf{p}_{1}d^{3}\mathbf{p}_{3}}{(2\pi)^{6}\sqrt{3}}(2\pi)^{3}\chi^{\mu}_{ss^{\prime}}$
$\langle L_{B},M_{L_{B}};S_{B},M_{S_{B}}|J_{B},M_{J_{B}}\rangle\delta^{3}(\mathbf{p}_{1}+\mathbf{p}_{3}-\mathbf{P}_{B})\;$
×$\times\psi_{L_{B}M_{L_{B}}}\left(\frac{m_{\bar{q_{i}}}\mathbf{p}_{1}-m_{q}\mathbf{p}_{3}}{m_{q}+m_{q_{i}}}\right)\;b^{\dagger}_{\mathbf{p}_{1}s_{1}}d^{\dagger}_{\mathbf{p}_{3}s_{3}}|0\rangle.$ (14)
The spatial wave function form of another final state meson $C$ is consistent with $B$. Here, $\psi_{{L}M_{L}}$ represent the spatial wave functions, which usually is taken as the simple harmonic oscillator wavefunction.
By combining the initial and final state wave functions with the Hamiltonian from Eq. (III), it is straightforward to obtain the matrix element $\langle BC|H_{I}|A\rangle=g\ (2\pi)^{3}\ \delta^{3}(\mathbf{P}_{A}-\mathbf{P}_{B}-\mathbf{P}_{C})\ M_{\ell J}(A\rightarrow BC)$ for the hybrid state $A$ decaying into mesons $B$ and $C$, which has been converted to partial wave amplitudes Ding:2006ya . The expression is as follows:
$M_{\ell J}$ $(A\rightarrow BC)=\sum_{\begin{array}[]{cccccc}M_{L_{g}},&\lambda_{g},&M_{L_{q\bar{q}}},&M_{S_{q\bar{q}}},&M_{L_{B}},&M_{S_{B}},\\ M_{L_{C}},&M_{S_{C}},&M_{J_{B}},&M_{J_{C}},&M_{J},&M_{\ell}\end{array}}{\cal C}\;{\cal F}{\cal S}(M_{S_{q\bar{q}}},\lambda_{g},M_{S_{B}},M_{S_{C}})I(M_{L_{q\bar{q}}},M_{L_{g}},M_{L_{B}},M_{L_{C}},M_{\ell})$ (17)
$\langle L_{g},M_{L_{g}};1,\lambda_{g}|J_{g},M_{L_{g}}+\lambda_{g}\rangle\langle L_{q\bar{q}},M_{L_{q\bar{q}}};J_{g},M_{L_{g}}+\lambda_{g}|L,M_{L_{g}}+\lambda_{g}+M_{L_{q\bar{q}}}\rangle\langle L,M_{L_{g}}+\lambda_{g}+M_{L_{q\bar{q}}};S_{q\bar{q}},M_{S_{q\bar{q}}}|J_{A},M_{J_{A}}\rangle$
$\langle L_{B},M_{L_{B}};S_{B},M_{S_{B}}|J_{B},M_{J_{B}}\rangle\langle L_{C},M_{L_{C}};S_{C},M_{S_{C}}|J_{C},M_{J_{C}}\rangle\langle J_{B},M_{J_{B}};J_{C},M_{J_{C}}|J,M_{J}\rangle\langle\ell,M_{\ell};J,M_{J}|J_{A},M_{J_{A}}\rangle.$ (18)
Here, ${\cal C}=\frac{2}{3}$ is the color overlap factor, ${\cal F}$ is the flavor overlap factor, ${\cal S}$ is the spin overlap factors and $I(M_{L_{q\bar{q}}},M_{L_{g}},M_{L_{B}},M_{L_{C}},M_{\ell})$ is the spatial overlap factor.
The spin overlap factor is:
${\cal S}$ $(M_{S_{q\bar{q}}},\lambda_{g},M_{S_{B}},M_{S_{C}})=\sum_{S}\sqrt{6(2S_{B}+1)(2S_{C}+1)(2S_{q\bar{q}}+1)}$
×$\times\left\{\begin{array}[]{lll}\frac{1}{2}&\frac{1}{2}&S_{B}\\ \frac{1}{2}&\frac{1}{2}&S_{C}\\ S_{q\bar{q}}&1&S\end{array}\right\}\langle S_{q\bar{q}},M_{S_{q\bar{q}}};1,\lambda_{g}|S,M_{S_{B}}+M_{S_{C}}\rangle$ (22)
$\langle S_{B},M_{S_{B}};S_{C},M_{S_{C}}|S,M_{S_{B}}+M_{S_{C}}\rangle.$ (23)
Eq. (III) indicates a decay selection rule, the “spin selection rule”: if the $q\bar{q}$ in hybrid is in a net spin single configuration then decay into final states consistent only of spin single states is forbidden, which also proposed in Ref. Page:1998gz . The flavor overlap factor ${\cal F}$ is:
${\cal F}=\sqrt{(2I_{B}+1)(2I_{C}+1)(2I_{A}+1)}\left\{\begin{array}[]{lll}i_{1}&i_{3}&I_{B}\\ i_{2}&i_{4}&I_{C}\\ I_{A}&0&I_{A}\end{array}\right\}\eta\ \varepsilon,$ (27)
where I (i) represents the isospin quantum number corresponding to the hadron (quark), $\eta=1$ if the gluon goes into strange quarks and $\eta=\sqrt{2}$ if gluon goes into non-strange ones, which means the ability of gluon to annihilate into non-strange quarks is stronger than that of annihilating into strange quarks. $\varepsilon$ is the number of the diagrams contributing to the decay. Finally the spatial overlap is given by:
$I($ $M_{L_{q\bar{q}}},M_{L_{g}},M_{L_{B}},M_{L_{C}},M_{\ell})=\int\frac{d^{3}\mathbf{p}d^{3}\mathbf{k}}{\sqrt{2\omega_{g}}(2\pi)^{6}}\;$
×$\times\psi_{L_{q\bar{q}}M_{L_{q\bar{q}}}}(\mathbf{P}_{B}-\mathbf{p})\psi_{L_{g}M_{L_{g}}}(\mathbf{k})\;\psi^{*}_{L_{B}M_{L_{B}}}\left(\frac{m_{\bar{q_{i}}}\mathbf{P}_{B}}{m_{q}+m_{\bar{q_{i}}}}-\mathbf{p}-\frac{\mathbf{k}}{2}\right)$
×$\times\psi^{*}_{L_{C}M_{L_{C}}}\left(-\frac{m_{q_{i}}\mathbf{P}_{B}}{m_{q_{i}}+m_{\bar{q}}}+\mathbf{p}-\frac{\mathbf{k}}{2}\right)\;d\Omega_{B}{Y^{M_{\ell}}_{\ell}}^{*}(\Omega_{B}).$ (28)
The calculation results indicate that the presence of the spatial integral term leads to the decay width of the excited state being lower than that of ground state.
The partial decay width is Ding:2006ya :
$\Gamma_{\ell J}(A\rightarrow BC)=\frac{\alpha_{s}}{\pi}\frac{p_{B}E_{B}E_{C}}{M_{A}}|M_{\ell J}(A\rightarrow BC)|^{2}.$ (29)
where $\alpha_{s}\approx 0.7\pm 0.3$, which represents the infrared quark-gluon vertex coupling Luna:2006tw ; Natale:2006nv ; Aguilar:2001zy .
In our calculations, the parameters are set as follows Ding:2006ya : $\beta_{q\bar{q}}=0.3\rm{GeV}$, $m_{s}=0.55\rm{GeV}$, $m_{u}=m_{d}=0.33\rm{GeV}$, $\omega_{g}=0.8\rm{GeV}$, and $\beta_{B}=\beta_{C}=\beta_{g}=0.4\rm{GeV}$.
Selection rules play a significant role in hybrid decays. For instance, the spin selection rule prohibits decay into final states composed entirely of spin-singlet mesons if the hybrid’s $q\bar{q}$ configuration is in a spin-singlet state. Additionally, spatial overlap integrals, which describe the overlap of wavefunctions between initial and final states, generally suppress the decay widths of excited states compared to ground states.
Another decay selection rule indicates that, assuming $\beta_{B}=\beta_{C}$, the decay of the GE state into $S+S$ is forbidden Iddir:2000yb . This rule arises from the treatment of integrals under the condition $\beta_{B}=\beta_{C}$. In Ref. Iddir:1988jd , the integrals are evaluated with $\beta_{B}\neq\beta_{C}$, and the results indicate that the spatial overlap is generally proportional to $(\beta_{B}-\beta_{C})^{2}$, which is typically small. The same outcome is observed in the flux tube model Close:1994hc . Similarly, if the two S-wave states possess different internal structures or sizes, as discussed in Ref. Close:1994pr , the selection rule prohibiting $S+S$ as final states no longer applies.
IV RESULTS AND ANALYSIS
IV.1 The $I^{G}J^{PC}=1^{-}1^{-+}$ state
The combined analysis of masses and decay widths provides essential insights into the nature of exotic mesons, particularly those with $J^{PC}=1^{-+}$. Section II presents mass calculations for different flavor configurations, showing that the $q\bar{q}g$, $(u\bar{u}-d\bar{d})g/\sqrt{2}$, and $(u\bar{u}+d\bar{d})g/\sqrt{2}$ states are mass-degenerate. Given the quantum numbers $I^{G}J^{PC}=1^{-}1^{-+}$, we assume that $\pi_{1}(1600)$ corresponds to the configuration $(u\bar{u}-d\bar{d})g/\sqrt{2}$ in our calculations. For the decay channels involving $f_{1}(1285)$ and $f_{1}(1420)$, we model them as follows:
$\begin{pmatrix}|f_{1}(1285)\rangle\\ |f_{1}(1420)\rangle\end{pmatrix}$ $=\begin{pmatrix}\cos\alpha&-\sin\alpha\\ \sin\alpha&\cos\alpha\end{pmatrix}\begin{pmatrix}|n\bar{n}\rangle\\ |s\bar{s}\rangle\end{pmatrix},$ (30)
where $\alpha\approx 30^{\circ}$ Dudek:2013yja .
Viewing $\pi_{1}(1600)$ as a hybrid state, we calculate its total decay width to be 15.15 MeV, with $b_{1}(1235)\pi$ identified as the primary decay channel (see Table 3). Reference Page:1998gz also reports a similar narrow decay width. However, this theoretical result exhibits a substantial deviation from the experimental values and averages reported by the PDG ParticleDataGroup:2024cfk . In our analysis, the decay width to $b_{1}(1235)\pi$ via $D$-wave is 0 MeV, implying that $\pi_{1}(1600)$ can only decay into $b_{1}(1235)\pi$ through $S$-wave. This finding contradicts the PDG’s reported branching ratio $\mathcal{B}\left(\left(b_{1}\pi\right)_{\text{D-wave}}\right)/\mathcal{B}\left(\left(b_{1}\pi\right)_{\text{S-wave}}\right)=0.3\pm 0.1$ ParticleDataGroup:2024cfk . This discrepancy raises questions about its classification as a hybrid state and points to alternative interpretations, such as tetraquark or molecular configurations.
Table 3: Partial and total decay widths of the $\pi_{1}(1600)$ state (in MeV ×$\times\alpha_{s}$). The final states include all charge conjugate pairs.
$b_{1}(1235)\pi$ $f_{1}(1285)\pi$ Total Experimental ParticleDataGroup:2024cfk
14.1 1.05 15.15 $370^{+50}_{-60}$
Similarly, $\pi_{1}(2015)$, although its mass matches lattice QCD predictions for the first excited hybrid state, exhibits narrow decay widths inconsistent with experimental observations, suggesting it may belong to a different category of exotic mesons. With the same exotic quantum numbers $I^{G}J^{PC}=1^{-}1^{-+}$, the mass of $\pi_{1}(2015)$ is reported to be $2014\pm 20^{+16}_{-16}$ MeV ParticleDataGroup:2024cfk , which is in close proximity to 2193 MeV (see Table 1). As suggested by Meyer:2010ku and previous lattice QCD (LQCD) mass calculations Dudek:2010wm , this approximate mass motivates us to propose that the $\pi_{1}(2015)$ state could be the first excited state of the hybrid $1^{-}1^{-+}$. Consequently, we assume that the $\pi_{1}(2015)$ state has the flavor configuration $(u\bar{u}-d\bar{d})g/\sqrt{2}$ and proceed with the calculation of its decay width. For the decay channels involving $K_{1}(1270)K$ and $K_{1}(1400)K$, we treat them as linear combinations of ${}^{1}P_{1}$ and ${}^{3}P_{1}$ states Page:1998gz ; Kokoski:1985is ,
$\begin{array}[]{l}|K_{1}(1270)\rangle=\sqrt{\frac{2}{3}}|K_{1}(^{1}P_{1})\rangle+\sqrt{\frac{1}{3}}|K_{1}(^{3}P_{1})\rangle\\ |K_{1}(1400)\rangle=-\sqrt{\frac{1}{3}}|K_{1}(^{1}P_{1})\rangle+\sqrt{\frac{2}{3}}|K_{1}(^{3}P_{1})\rangle\end{array}$ (31)
and the decay channels associated with $h_{1}(1170)$, $h_{1}(1380)$, we regard them as
$\begin{pmatrix}|h_{1}(1170)\rangle\\ |h_{1}(1380)\rangle\end{pmatrix}$ $=\begin{pmatrix}\sin\alpha_{i}&\cos\alpha_{i}\\ \cos\alpha_{i}&-\sin\alpha_{i}\end{pmatrix}\begin{pmatrix}|n\bar{n}\rangle\\ |s\bar{s}\rangle\end{pmatrix},$ (32)
$\alpha_{i}=78.7^{\circ}$, see Ref. Wang:2019qyy .
As shown in Table 4, the total decay width of $\pi_{1}(2015)$ that we calculated is 0.475 MeV, which is inconsistent with the experimental values ParticleDataGroup:2024cfk . The primary decay channel of $\pi_{1}(2015)$ is $b_{1}(1235)\pi$, while other decay channels can be neglected. Experimentally, $\pi_{1}(2015)$ has been observed in the $b_{1}(1235)\pi$ channel. However, the narrow decay width leads us to question the interpretation of $\pi_{1}(2015)$ as the first excited state of the $1^{-}1^{-+}$ hybrid. Until now, $\pi_{1}(2015)$ has been observed only in the $b_{1}(1235)\pi$ and $f_{1}(1285)\pi$ channels E852:2004rfa ; E852:2004gpn .
Table 4: The partial and total widths of the strong decays of the $\pi_{1}(2015)$ state.(width is in MeV ×$\times\alpha_{s}$ for the channels)
$f_{1}(1285)\pi$ $h_{1}(1170)\rho$ $b_{1}(1235)\omega$ $b_{1}(1235)\pi$ $K_{1}(1270)K$
0.03 $\approx 0$ $\approx 0$ 0.42 0.015
$K_{1}(1400)K$ $a_{1}(1260)\rho$
$\approx 0$ 0.01
Total Exp. ParticleDataGroup:2024cfk
[…]
Furthermore, we predict the properties of the first excited isoscalar hybrids $\pi_{1}(2193)$ with $I^{G}J^{PC}=1^{-}1^{-+}$. Its decay properties are listed in Table 5. This state is expected to have narrow decay widths and distinctive dominant decay channels, such as $K_{1}(1650)K$, and $b_{1}(1235)\pi$. For the decay channel involving $K_{1}(1650)K$, we consider $K_{1}(1650)$ to be a linear combination of the $2^{1}P_{1}$ and $2^{3}P_{1}$ states, based on the mass predictions of mesons in Ref. Godfrey:1985xj , and assume it adopts the ideal mixing angle. Following the approach in Ref. Blundell:1996as , we assume a mixing angle of $45^{\circ}$ for $K_{1}(1650)^{+}$ and $K_{1}(1650)^{0}$, and $-45^{\circ}$ for $K_{1}(1650)^{-}$ and $\bar{K}_{1}(1650)^{0}$. The narrow width make these states challenging to observe experimentally, but their detection would provide valuable confirmation of the theoretical framework for hybrid mesons.
Table 5: The partial and total widths of the strong decays of the $\pi_{1}(2193)$ which is assumed as the first excited state of $1^{-}1^{-+}$ hybrid.(width is in MeV ×$\times\alpha_{s}$ for the channels)
$f_{1}(1285)\pi$ $K_{1}(1650)K$ $b_{1}(1235)\pi$
$0.13$ $5.15$ 1.36
$\rho(1450)\pi$ $\eta_{2}(1645)\pi$
$0.2$ $0.2$
Total Exp.
$7.07$-
IV.2 The ground states of $I^{G}J^{PC}=0^{+}1^{-+}$
As mentioned in Section II, the mass of the pure $s\bar{s}g$ state is 2023 MeV, which is higher than 1855 MeV, whereas the pure $(u\bar{u}+d\bar{d})g/\sqrt{2}$ state has a mass lower than 1855 MeV. This suggests that $\eta_{1}(1855)$ is likely a mixed state of $s\bar{s}g$ and $(u\bar{u}+d\bar{d})g/\sqrt{2}$, as proposed by Swanson:2023zlm . This hypothesis also lends support to efforts to construct the $1^{-+}$ nonet Chen:2023ukh ; Qiu:2022ktc ; Eshraim:2020ucw .
As proposed in Refs. Chen:2023ukh ; Eshraim:2020ucw ; Shastry:2023ths , we assume that $\eta_{1}(1855)$ is the higher state and $\eta_{1}(1600)$ is the lower state of the $1^{-+}$ hybrid isoscalar nonet. In this paper, we follow this assumption. The mixing can be parameterized as follows:
$\begin{pmatrix}|\eta_{1}^{H}\rangle\\ |\eta_{1}^{L}\rangle\end{pmatrix}$ $=\begin{pmatrix}\sin\theta&\cos\theta\\ \cos\theta&-\sin\theta\end{pmatrix}\begin{pmatrix}|n\bar{n}g\rangle\\ |s\bar{s}g\rangle\end{pmatrix},$ (33)
$\theta$ is the mixing angle and $n\bar{n}=(u\bar{u}+d\bar{d})/\sqrt{2}$.
Similar to the flavor mixing treatment of meson decay width in Ref. Li:2008mza , the partial widths of $\eta_{1}^{H}$ and $\eta_{1}^{L}$ become with mixing:
$\Gamma(\eta_{1}^{H}\rightarrow BC)=G\sum_{LS}|\sin\phi{\cal{M}}^{LS}_{n\bar{n}\rightarrow BC}+\cos\phi{\cal{M}}^{LS}_{s\bar{s}\rightarrow BC}|^{2},$
$\Gamma(\eta_{1}^{L}\rightarrow BC)=G^{{}^{\prime}}\sum_{LS}|\cos\phi{\cal{M}}^{LS}_{n\bar{n}\rightarrow BC}-\sin\phi{\cal{M}}^{LS}_{s\bar{s}\rightarrow BC}|^{2}.$
Where $G=\frac{\alpha_{s}PE_{B}E_{C}}{\pi M_{\eta_{1}^{H}}}$, $G^{{}^{\prime}}=\frac{\alpha_{s}PE_{B}E_{C}}{\pi M_{\eta_{1}^{L}}}$, for the decay channels associated with $\eta$ and $\eta^{\prime}$, we regard them as
$\begin{pmatrix}\eta\\ \eta^{\prime}\end{pmatrix}$ $=\begin{pmatrix}\cos\theta_{i}&-\sin\theta_{i}\\ \sin\theta_{i}&\cos\theta_{i}\end{pmatrix}\begin{pmatrix}|n\bar{n}\rangle\\ |s\bar{s}\rangle\end{pmatrix},$ (34)
$\theta_{i}\approx 43.3^{\circ}$, see Ref. Wang:2020due .
In contrast, assuming that $\eta_{1}(1855)$ is a higher isospin partner, it exhibits properties that strongly support its interpretation as a mixed hybrid state. Table 6 presents the decay results. Its dominant decay channels, $a_{1}(1260)\pi$ and $K_{1}(1270)K$ align well with theoretical predictions (see Fig. 3(a)), and the mixing angle between $(u\bar{u}+d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$ is constrained to the range of $21^{\circ}$ to $74^{\circ}$. This result presents a compelling argument for classifying $\eta_{1}(1855)$ as a hybrid meson and highlights the importance of mixing effects in understanding its decay dynamics.
For quantum numbers $I^{G}J^{PC}=0^{+}1^{-+}$, the state $(u\bar{u}+d\bar{d})g/\sqrt{2}$ is assigned as the lower isoscalar partner to the $\eta_{1}(1855)$ state, with the assumption that $\eta_{1}(1855)$ is the higher isoscalar partner. The lower state is labeled as $\eta_{1}(1600)$, with the mixing angle ranging from $21^{\circ}$ to $74^{\circ}$, as previously discussed. The results presented in Table 7 show that $a_{1}(1260)\pi$ is the characteristic decay channel. The total decay width of $\eta_{1}(1600)$, as a function of the mixing angle, is shown in Figure 3 (b). The decay width of the $\eta_{1}(1600)$ state is predicted to be narrow, consistent with results reported in Chen:2023ukh ; Shastry:2023ths .
The results highlight the need for precise experimental measurements of decay branching ratios, particularly for $\alpha_{s}$ of its decay channels will further constrain theoretical models and provide clarity on its hybrid nature. Similarly, targeted searches for the predicted excited states will be crucial in expanding our understanding of the $1^{-+}$ meson spectrum.
Table 6: The partial and total decay widths of the $\eta_{1}(1855)$, which was assumed as a higher state. $s\equiv\sin\theta,c\equiv\cos\theta$. (width is in MeV ×$\times\alpha_{s}$ for the channels)
$a_{1}(1260)\pi$ $f_{1}(1285)\eta$ $\pi(1300)\pi$
$58.2s^{2}$ $1.2c^{2}+2s^{2}+3.1cs$ $3.48s^{2}$
$K_{1}(1270)K$
$106.6c^{2}+58.8s^{2}+158.3cs$
Total Exp. ParticleDataGroup:2024cfk
$107.8c^{2}+122.5s^{2}+161.4cs$ $188\pm 18^{+3}_{-8}$
Figure 3: The decay width of $\eta_{1}(1855)$ and $\eta_{1}(1600)$ which was assumed as a higher state and lower state, respectively. (a) represents total and partial decay of $\eta_{1}(1855)$, (b) represents total and partial decay of $\eta_{1}(1600)$. The yellow shaded area represents the range of the experimental decay width. The blue shaded area denotes the range of theoretical mixing angles.
Table 7: The partial and total decay widths of the $\eta_{1}(1600)$, which was assumed as a lower state. $s\equiv\sin\theta,c\equiv\cos\theta$. (width is in MeV ×$\times\alpha_{s}$ for the channels)
$a_{1}(1260)\pi$ Total Exp.
$28.6c^{2}$ $28.6c^{2}$-
IV.3 The first excited states of $I^{G}J^{PC}=0^{+}1^{-+}$
In accordance with the mass spectrum calculated in Section II, the first excitation state of $\eta_{1}(1855)$ is labeled as $\eta_{1}(2355)$, maintaining the same mixing angle as that of $\eta_{1}(1855)$. The decay width results are presented in Table 8. For a clearer visualization of the decay widths, refer to Fig. 4(a). The total decay width of $\eta_{1}(2355)$ is approximately 100 MeV, with dominant decay channels including $K_{1}(1270)K$, $K_{1}(1650)K$, and $a_{1}(1640)\pi$.
Additionally, the first excited state of $\eta_{1}(1600)$ is named $\eta_{1}(2193)$, serving as the partner of $\eta_{1}(2355)$. The decay widths are presented in Table 9. As shown in Fig. 4 (b), the main decay modes of $\eta_{1}(2193)$ include $K_{1}(1270)K$, $K_{1}(1650)K$, and $a_{1}(1640)\pi$. We anticipate that future experiments will investigate these channels to detect $\eta_{1}(2355)$ and $\eta_{1}(2193)$.
Table 8: The partial and total widths of the strong decays of the $\eta_{1}(2355)$ , which was assumed as a higher state. $s\equiv\sin\theta,c\equiv\cos\theta$. (width is in MeV ×$\times\alpha_{s}$ for the channels)
$b_{1}(1235)\rho$ $a_{1}(1260)\pi$ $f_{1}(1420)\eta$
$1.08s^{2}$ $5.64s^{2}$ $\approx 0$
$K_{1}(1270)K$ $K_{1}(1400)K$ $K_{1}(1270)K^{*}$
$1s^{2}+17.6c^{2}+8.4cs$ $\approx 0$ $\approx 0$
$K_{1}(1650)K$ $\pi(1300)\pi$ $\pi_{2}(1670)\pi$
$22.84s^{2}+36.2c^{2}+57.5cs$ $6.3s^{2}$ $7.8s^{2}$
$K(1460)K$ $K^{*}(1410)K$ $a_{1}(1640)\pi$
$0.5s^{2}+3.36c^{2}+2.53cs$ $0.31s^{2}+2.08c^{2}+1.6cs$ $35.28s^{2}$
Total Exp.
$80.75s^{2}+59.24c^{2}+70.03cs$-
Figure 4: The decay width of $\eta_{1}(2355)$ and $\eta_{1}(2193)$ predicted by this work. (a) represents total and partial decay of $\eta_{1}(2355)$ , (b) represents total and partial decay of $\eta_{1}(2193)$. The shaded area denotes the range of theoretical mixing angles.
Table 9: The partial and total widths of the strong decays of the $\eta_{1}(2193)$ ,which was assumed as a lower state. $s\equiv\sin\theta,c\equiv\cos\theta$. (width is in MeV ×$\times\alpha_{s}$ for the channels)
$a_{1}(1260)\pi$ $a_{1}(1640)\pi$ $K_{1}(1270)K$
$2.7c^{2}$ $26c^{2}$ $7.52s^{2}+0.12c^{2}-1.89cs$
$\pi(1300)\pi$ $\pi_{2}(1670)\pi$ $K(1460)K$
$3.6c^{2}$ $2.82c^{2}$ $1.1s^{2}+0.16c^{2}-0.84cs$
$K_{1}(1400)K$ $K_{1}(1650)K$
$\approx 0$ $19s^{2}+10c^{2}-27.5cs$
Total Exp.
$27.62s^{2}+45.4c^{2}-30.23cs$-
V Discussion and conclusion
This study investigates the properties of hybrid $J^{PC}=1^{-+}$ mesons and their isoscalar counterparts through comprehensive mass and decay analyzes. Using a potential model inspired by the lattice QCD and the constituent-gluon framework, we calculate the masses and decay widths of key exotic mesons, shedding light on their structure and classification.
Our findings suggest that $\eta_{1}(1855)$ is a strong candidate for a mixed hybrid meson, composed of $(u\bar{u}+d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$ components, with mixing angles constrained between $21^{\circ}$ to $74^{\circ}$. Its mass and decay properties, including dominant channels such as $a_{1}(1260)\pi$ and $K_{1}(1270)K$ , align well with theoretical predictions. These results support its classification as a hybrid meson within the meson nonet.
In contrast, the narrow decay widths predicted for $\pi_{1}(1600)$ and $\pi_{1}(2015)$ deviate significantly from experimental observations, challenging their classification as hybrid mesons. Alternative explanations, such as tetraquark or molecular configurations, merit further investigation. Our results also predict narrow width excited states, $\eta_{1}(2355)$ and $\eta_{1}(2193)$, which primarily decay into channels like $K_{1}(1270)K$ and $a_{1}(1640)\pi$. While challenging to detect experimentally, these states offer promising opportunities to test and validate hybrid meson models.
This work underscores the importance of precise experimental measurements to resolve ambiguities in hybrid meson classification. Future experiments should focus on measuring branching ratios for key decay channels, such as those of $\eta_{1}(1855)$, to refine mixing angle estimates. Furthermore, searches for the predicted excited states will be instrumental in advancing our understanding of hybrid meson dynamics and their role in exotic hadronic states.
The results presented here highlight the intricate interplay between theory and experiment in the study of exotic mesons. Ongoing collaboration between experimental efforts and theoretical advancements, including refined lattice QCD calculations and improved decay models, will be critical in uncovering the true nature of hybrid mesons and enhancing our understanding of the strong interaction.
Acknowledgements
This work is supported by the National Natural Science Foundation of China under Grant No. 12405104, the Natural Science Foundation of Hebei Province under Grant No. A2022203026, the Higher Education Science and Technology Program of Hebei Province under Contract No. BJK2024176, and the Research and Cultivation Project of Yanshan University under Contract No. 2023LGQN010.